OpenStudy (anonymous):

make L the subject Z=√R^2+(2πFL-1/2πFC)^2

OpenStudy (sasogeek):

if i'm reading this right, tell me if this is the equation \[\huge z=\sqrt{R^2}+(\large \frac{2\pi FL-1}{2\pi FC}\huge ) \] is this it?

OpenStudy (sasogeek):

and the bracket is raised to the power 2...

OpenStudy (angela210793):

or this one?